# OpenDosimeter: Open Hardware Personal X-ray Dosimeter

### v1

@@ Abstract
We present OpenDosimeter (www.opendosimeter.org), an open hardware solution for real-time personal X-ray dose monitoring based on a scintillation counter. Using an X-ray sensor assembly (LYSO + SiPM) on a custom board powered by a Raspberry Pi Pico, OpenDosimeter provides real-time feedback (1 Hz), data logging (10 hours), and battery-powered operation. One of the core innovations is that we calibrate the device using 241Am found in ionization smoke detectors. Specifically, we use the $\gamma$-emissions to spectrally calibrate the dosimeter, then calculate the effective dose from X-ray exposure by compensating for the scintillator absorption efficiency and applying energy-to-dose coefficients derived from tabulated data in the ICRP 116 publication. We demonstrate that this transparent approach enables real-time dose rate readings with a linear response between 0.1–1000 µSv/h at $\pm$25accuracy, tested for energies up to 120 keV. The maximum dose rate readings are limited by pile-up effects when approaching count rate saturation ($\sim$77 kcps at $\sim$13 µs average pulse processing time). The total component cost for making an OpenDosimeter is <\$100, which, combined with its open design (both hardware and software), enables cost-effective local reproducibility on a global scale. This paper complements the open-source documentation by explaining the underlying technology, the algorithm for dose calculation, and areas for future improvement.
@@ Introduction
Monitoring occupational X-ray exposure is critical for ensuring the short- and long-term safety of radiation workers. To minimize health risks, regulations set dose limits for these workers, such as a yearly effective dose below 20 mSv (averaged over five consecutive years) and not exceeding 50 mSv in any single year [1]. Yet, while there is limited data on the global availability, many workers in low-resource settings have little to no access to personal dosimeters. Traditional dosimeters can be prohibitively expensive and require complicated logistics around calibration, readout, and reporting. All of this leads to radiation workers having restricted insight into their own exposure levels.
Personal dosimeters are categorized as either active or passive devices. On the passive side, dosimeters based on thermoluminescence (TLDs) or optically stimulated luminescence (OSL) allow for cost-effective scalability and are often offered as a subscription service by an external party. However, individuals wearing these dosimeters receive infrequent feedback on their radiation exposure, typically only monthly or quarterly. Additionally, the logistics of collecting dosimeters and sending them to remote facilities for readout are challenging. Moreover, the lack of immediate feedback on exposure levels hinders the individual in developing and maintaining effective radiation safety practices.
On the active side, there are real-time electronic personal dosimeters. An example is the RaySafe i3 marketed towards interventional radiology procedures where dose exposure to the operators can be high which makes direct feedback critical. Unfortunately, these active devices are often prohibitively expensive (>\$1000) for scaling to larger groups of radiation workers.
@@
We present OpenDosimeter (www.opendosimeter.org), an open hardware solution for low-cost and real-time personal radiation monitoring. This manuscript complements the open-source documentation (see our GitHub repository) by explaining the underlying technology, the algorithms and procedures for spectral calibration and dose calculation. Lastly, we benchmark its performance against a commercially available active dosimeter (RaySafe i3).
@@ Results
Figure 2: Calibration and dose calculation. a, Disassembled ionization smoke detector with 241Am used for spectral calibration b, Raw calibration histogram (light purple; shaded) and moving average (dark purple; solid) showing two peaks. Inset: zoom on the two peaks with the underlying 241Am $\gamma$-spectrum overlaid (measured with a CdTe spectrometer). The spectral calibration performs a linear fit of the two peaks with the expected 241Am emission energies. c, LYSO (1 mm thick) spectral absorption efficiency. d, Per-photon-dose spectral coefficients, derived from ICRP 116 (Table A.1; AP direction) and scaled with the LYSO crystal area (25 mm2). e, Algorithm for calculating the dose rate and accumulated dose.
@@
The complete OpenDosimeter is housed in a 3D-printed case measuring 73 $\times$ 42 $\times$ 23 mm (Fig. 1f), with its compact size making it wearable on the chest or belt for monitoring personal dose from occupational X-ray exposure.
The device firmware implements a simple state machine for user interaction (Fig. 1g). Upon power-on, the device enters the main operational mode. Short button presses cycle through displaying the current dose rate and the accumulated dose. The options menu is entered through a long button press, where a short press resets the dose values, and a long press brings the device to the calibration procedure (see next section). Dose values are continuously logged at 1 Hz with values for the past 10 hours can be accessed and visualized through our Web Interface. The interface does not require any software installation on the user end, and can be accessed simply by connecting the OpenDosimeter over USB to any computer with an internet connection. The USB connection is also used to charge the lithium-ion polymer (LiPO) battery. A battery with 1200 mAh capacity currently enables up to 20 hours of operation (power consumption typically <70 mA @ 3.7V).
@@
First, we evaluated the photon count rate performance of the OpenDosimeter (Fig. 3a). The output count rate increases linearly with the input rate up to approximately 100 kcps (thousand counts per second), defining a reliable photon counting region (shaded purple). Beyond this point, the output becomes unreliable likely due to increasing probability of pulse pile-up effects. The maximum output count rate plateaus at $\sim$77 kcps, corresponding to a deadtime fraction approaching 100%. The latter is estimated within the OpenDosimeter by multiplying the output count rate with the average per-pulse sampling time (cf. Fig. 1d, $\sim$13 µs) which ultimately defines the maximum output count rate. Within the reliable photon counting region, the estimated deadtime percentage can be used as a correction factor for the output count rate.
@@
Next, we demonstrate the real-time performance of the OpenDosimeter in a dynamic radiation environment compared with the commercial dosimeter (Fig. 3d). The dosimeters were simultaneously exposed to X-rays at increasing tube currents, indicated by multiplicative factors noted as horizontal lines in the graph. Each exposure lasted 5 seconds. The plot shows both the instantaneous dose rate and the accumulated dose as measured by both devices. OpenDosimeter closely tracks the commercial device in both dose rate response and accumulated dose, showcasing the ability to accurately monitor rapidly changing radiation levels. Lastly, the angular response of the OpenDosimeter is shown in Fig. 3e with data from the RaySafe i3 datasheet overlaid for comparison.
@@ Acknowledgements.
The OpenDosimeter project is a derivative of the Open Gamma Detector project [5]. We acknowledge funding support from the Knut and Alice Wallenberg Foundation, and the King Center for Global Development at Stanford University.
@@
N.G. and K.S. conceived the project. N.G., A.K., A.S.W., and K.S. designed the device and its functionality. G.A. and E.A. gave input on functionality and user design. M.R. and K.S. developed the hardware and software. N.R.B., J.W., and K.S. performed the characterization. K.S. analyzed the data. J.L. and K.S. wrote the open-source documentation. K.S. led the project and wrote the manuscript with input from all authors.

### v2

@@ Abstract
We present OpenDosimeter (www.opendosimeter.org), an open hardware solution for real-time personal X-ray dose monitoring based on a scintillation counter. Using an X-ray sensor assembly (LYSO + SiPM) on a custom board powered by a Raspberry Pi Pico, OpenDosimeter provides real-time feedback (1 Hz), data logging (10 hours), and battery-powered operation. One of the core innovations is that we calibrate the device using 241Am found in ionization smoke detectors. Specifically, we use the $\gamma$-emissions to spectrally calibrate the dosimeter, then calculate the effective dose from X-ray exposure by compensating for the scintillator absorption efficiency and applying energy-to-dose coefficients derived from tabulated data in the ICRP 116 publication. We demonstrate that this transparent approach enables real-time dose rate readings with a linear response between 0.1–1000 µSv/h at $\pm$25accuracy, tested for energies up to 120 keV. The maximum dose rate readings are limited by pile-up effects when approaching count rate saturation ($\sim$77 kcps at $\sim$13 µs average pulse processing time). The total component cost for making an OpenDosimeter is <\$100, which, combined with its open design (both hardware and software), enables cost-effective local reproducibility on a global scale. Through a student workshop, we also demonstrate its effectiveness as an educational and capacity-building tool. This paper complements the open-source documentation by explaining the underlying technology, the algorithm for dose calculation, and areas for future improvement.
@@ Introduction
Monitoring occupational X-ray exposure is critical for ensuring the short- and long-term safety of radiation workers. To minimize health risks, regulations set dose limits for these workers, such as a yearly effective dose below 20 mSv (averaged over five consecutive years) and not exceeding 50 mSv in any single year [1]. However, while there is limited data on the global availability, many workers in low-resource settings have little or no access to personal dosimeters. Traditional dosimeters can be prohibitively expensive and require complicated logistics around calibration, readout, and reporting. All of this leads to radiation workers having restricted insight into their own exposure levels.
Personal dosimeters are categorized as either active or passive devices. On the passive side, dosimeters based on thermoluminescence (TLDs) or optically stimulated luminescence (OSL) allow for cost-effective scalability and are often offered as a subscription service by an external party. However, individuals wearing these dosimeters receive infrequent feedback on their radiation exposure, typically only monthly or quarterly. In addition, the logistics of collecting dosimeters and sending them to remote facilities for reading is challenging. Moreover, the lack of immediate feedback hinders the individual in maintaining effective radiation safety practices.
On the active side, there are real-time electronic personal dosimeters. An example is the RaySafe i3 marketed towards interventional radiology procedures where dose exposure to the operators can be high, which makes direct feedback critical. Unfortunately, these active devices are often prohibitively expensive (>\$1000) for scaling to larger groups of radiation workers.
@@
We present OpenDosimeter (www.opendosimeter.org), an open hardware solution for low-cost and real-time personal radiation monitoring. This manuscript complements the open-source documentation (see our GitHub repository) by explaining the underlying technology, the algorithms for spectral calibration and dose calculation, as well as benchmarking against a commercial active dosimeter (RaySafe i3). Lastly, we demonstrate its potential for capacity-building through a student workshop.
@@ Results
Figure 2: Calibration and dose calculation. a, Disassembled ionization smoke detector with 241Am used for spectral calibration b, Raw calibration histogram (light purple; shaded) and moving average (dark purple; solid) showing two peaks. Inset: zoom on the two peaks with the underlying 241Am $\gamma$-spectrum overlaid (measured with a CdTe spectrometer). The spectral calibration performs a linear fit of the two peaks with the expected 241Am emission energies. c, LYSO (1 mm thick) spectral X-ray absorption efficiency. d, Per-photon-dose spectral coefficients, derived from ICRP 116 (Table A.1; AP direction) and scaled with the LYSO crystal area (25 mm2). e, Algorithm for calculating the dose rate and accumulated dose.
@@
The complete OpenDosimeter is housed in a 3D-printed case measuring 73 $\times$ 42 $\times$ 23 mm (Fig. 1f), with its compact size making it wearable on the chest or belt for monitoring dose from occupational X-ray exposure.
The device firmware implements a simple state machine for user interaction (Fig. 1g). Upon power-on, the device enters the main operational mode. Short button presses cycle through two display states; showing either the current dose rate or the accumulated dose. The options menu is entered through a long button press, where a short press resets the dose values, and a long press brings the device to the calibration procedure (see next section). Dose values are continuously logged at 1 Hz with values for the past 10 hours can be accessed and visualized through our Web Interface. The interface does not require any software installation on the user end, and can be accessed simply by connecting the OpenDosimeter over USB to any computer with an internet connection. The USB connection is also used to charge the lithium-ion polymer (LiPO) battery. A battery with 1200 mAh capacity currently enables up to 20 hours of operation per full charge (power consumption typically <70 mA @ 3.7V).
@@
First, we evaluated the photon count rate performance of the OpenDosimeter (Fig. 3a). The output count rate increases linearly with the input rate up to approximately 100 kcps (thousand counts per second), defining a reliable photon-counting region (shaded purple). Beyond this point, the output becomes unreliable likely due to increasing probability of pulse pile-up effects. The maximum output count rate plateaus at $\sim$77 kcps, corresponding to a deadtime fraction approaching 100%. The latter is estimated within the OpenDosimeter by multiplying the output count rate with the average per-pulse sampling time (cf. Fig. 1d, $\sim$13 µs) which ultimately defines the maximum output count rate. Within the reliable photon counting region, the estimated deadtime percentage can be used as a correction factor for the output count rate.
@@
Next, we demonstrate the real-time performance of the OpenDosimeter in a dynamic radiation environment compared with the commercial dosimeter (Fig. 3d). The dosimeters were simultaneously exposed to X-rays at increasing tube currents, indicated by multiplicative factors noted as horizontal lines in the graph. Each exposure lasted 5 seconds. The plot shows both the instantaneous dose rate and the accumulated dose as measured by both devices. OpenDosimeter closely tracks the commercial device in both dose rate response and accumulated dose, showcasing the ability to accurately monitor rapidly changing radiation levels. Lastly, the angular response of the OpenDosimeter is shown in Fig. 3e with the RaySafe i3 (from datasheet) overlaid for comparison.
Pilot workshop for capacity-building
Figure 4: OpenDosimeter as a tool for capacity-building in radiation technology. a, Workshop format spanning seven days, integrating stakeholder visits, learning activities, and design improvements by the student participants. b, Pre- and post-workshop survey results (n=20 and n=12, respectively) showing improvements in knowledge, confidence, and practical experience, scored from 1 (low) to 5 (high). c, Excerpt of reflections from entry and exit surveys highlighting participant motivations and learning outcomes (quotes lightly edited for brevity while preserving language and content). d, Our workshop as a case study on open hardware implementation towards locally sustainable innovation.
Beyond utility as a radiation monitoring device, we evaluated OpenDosimeter as a tool for education and capacity building through a pilot workshop at Strathmore University in Nairobi, Kenya (Fig. 4a). We enrolled local engineering students with minimal prior knowledge of X-ray radiation and provided them with components to build OpenDosimeters using the open-source documentation. We started the workshop with lectures on radiation safety and detection basics, then progressed to hands-on activities building dosimeters including hardware and software. We also invited local stakeholders from policy, metrology, and radiology domains to interact with the students and share their perspectives on accessible radiation monitoring.
We collected survey results at the beginning ("entry") and end ("exit") of the workshop to gain insight into the experiences of our workshop participants. The quantitative self-assessment surveys showed that by building an OpenDosimeter, our participants gained both theoretical understanding and practical capability to work with radiation technology (Fig. 4b). While most participants entered with minimal experience with both radiation detection and open hardware development, all workshop completers (n = 12, with some of the initial n = 20 participants dropping out) successfully assembled functioning OpenDosimeters. In their qualitative reflections (Fig. 4c), the students demonstrated enhanced understanding of both technical concepts and broader societal implications of accessible detection technologies for radiation safety. We found the combination of theoretical learning through lectures, hands-on experience building the OpenDosimeter hardware followed by software programming, effectively built capacity in radiation detection and safety among our participants.
As illustrated in Fig. 4a, our participants progressed from minimal initial knowledge to achieving technical understanding sufficient for independent innovation. In the final days of the workshop, participants identified and addressed a limitation in the initial software; the need to subtract intrinsic LYSO crystal background radioactivity from the dose rate calculations. Working collaboratively, the participants implemented this improvement and we published it as a new firmware version (v1.0.1).
On a general level, our workshop can be seen as a case study demonstrating a framework for using open hardware to build local technical capacity (Fig. 4d). The engagement of local stakeholders from clinical radiology, metrology, and regulatory policy helped our student participants recognize that there exists strong incentives for local implementation. These stakeholders emphasized how locally designed and manufactured dosimeters could bridge the accessibility gap in personal radiation monitoring among radiation workers locally, showing the participants how their practical knowledge in dosimeter technology could address these needs. The alignment between open hardware and local stakeholder interests paves the way for sustainable innovation, enabling young innovators to adapt open blueprints into locally relevant solutions. Further, the technical achievement of the software upgrade developed by the workshop participants (see earlier) shows how open hardware approaches can simultaneously address technological needs while building local innovation capacity, a combination that proprietary "black-box" hardware prevents by restricting deep technical understanding and modifications.
Our OpenDosimeter workshop thus demonstrates how combining open hardware with local stakeholder incentives creates a framework for developing adapted technical solutions while empowering next-generation innovators. The potential of this framework is perhaps best captured by one engineering student reflecting in the exit survey (Fig. 4c): "I have realized that people actually need these tools. And it’s crazy because I can make them."
@@ Discussion
Through our pilot workshop (Fig. 4) we demonstrate that open hardware can serve as an effective catalyst for sustainable local innovation when aligned with local stakeholder interests. By lowering barriers to entry through openly shared designs and documentation, we believe that projects like OpenDosimeter enable next-generation innovators to develop deep domain-specific technical competency. We propose that this approach to capacity building, combining technical education using open hardware with stakeholder incentives, could serve as a model for similar initiatives in other domains where accessible technology solutions are needed to address critical community needs.
@@
Pilot workshop for capacity-building
We conducted the pilot workshop (Fig. 4) over seven days at Strathmore University, Nairobi, Kenya between the dates of Aug 28 to Sept 6, 2024. We selected twenty undergraduate engineering students (50/50 female/male ratio) initially, with twelve completing the full workshop (the remaining dropped out throughout the workshop period due to courses and other obligations). The workshop was held physically at the Makerspace Research Lab at the School of Computing and Engineering Sciences, which had the equipment needed to build the OpenDosimeter hardware (soldering stations, oscilloscopes, and other basic electronic tools). The student participants performed software programming of the devices using their own laptops.
On the stakeholder side, we engaged with a director at the Kenya Nuclear Regulatory Authority (regulatory policy) who visited the makerspace during the workshop. The workshop curriculum also included visits to the Mama Lucy Kibaki Hospital (clinical radiology) as well as the Secondary Standards Dosimetry Laboratory at the Kenya Bureau of Standards (metrology).
@@ Acknowledgements.
The OpenDosimeter project is a derivative of the Open Gamma Detector project [5]. We thank Shirin Pourashraf at Stanford University for assisting with SiPM configurations. We also thank the students who participated in the workshop at Strathmore University for their feedback on the OpenDosimeter design and its reproducibility. We acknowledge funding support from the Knut and Alice Wallenberg Foundation, and the King Center for Global Development at Stanford University.
@@
N.G. and K.S. conceived the project. N.G., A.K., A.S.W., and K.S. designed the device and its functionality. G.A. and E.A. gave input on functionality and user design. M.R. and K.S. developed the hardware and software. N.R.B., J.W., and K.S. performed the characterization. P.K. and K.S. organized the workshop at Strathmore University. N.R.B., P.K., and K.S. ran the workshop. K.S. analyzed the data. J.L. and K.S. wrote the open-source documentation. K.S. led the project and wrote the manuscript with input from all authors.